A large rectangular movie screen in an Imax theater as an area of 10,384 square feet. Find the dimensions of the screen if it is 30 feet longer than it is wide?
Answer by geetha_rama(94)

Let x be the width of the screen and
x + 30 be the length of the screen
Area of the screen = 10384 square feet
=>x(x+30) = 10384
x^2 + 30x = 10384
x^2 + 30x - 10384 = 0
x2+118x - 88x -10384 = 0
x(x+118) -88(x+118) = 0
(x+118)(x-88) = 0
x=-118,88
Since width should be positive,
width = 88 feet
length = 118 feet